# Separating Tongue from Thought: Activation Patching Reveals Language-Agnostic Concept Representations in Transformers

### v3

@@ Abstract
A central question in multilingual language modeling is whether large language models (LLMs) develop a universal concept representation, disentangled from specific languages. In this paper, we address this question by analyzing latent representations (latents) during a word-translation task in transformer-based LLMs. We strategically extract latents from a source translation prompt and insert them into the forward pass on a target translation prompt. By doing so, we find that the output language is encoded in the latent at an earlier layer than the concept to be translated. Building on this insight, we conduct two key experiments. First, we demonstrate that we can change the concept without changing the language and vice versa through activation patching alone. Second, we show that patching with the mean representation of a concept across different languages does not affect the models’ ability to translate it, but instead improves it. Finally, we generalize to multi-token generation and demonstrate that the model can generate natural language description of those mean representations. Our results provide evidence for the existence of language-agnostic concept representations within the investigated models. [^1] Code for reproducing our experiments is available at https://anonymous.4open.science/r/llm-lang-agnostic-1A75/
@@
Clément Dumas* 1,2  Chris Wendler* 3  Veniamin Veselovsky3  Giovanni Monea3,4  Robert West3 1ENS Paris-Saclay  2Université Paris-Saclay  3EPFL Lausanne  4Cornell Tech *Equal contribution {clement.dumas@ens-paris-saclay.fr, chris.wendler@epfl.ch}
@@ 1 Introduction
l question in multilingual language modeling is whether LLMs develop universal concept representations that transcend specific languages Wendler et al. (2024). For example, when a model processes the word "cat" in English and "chat" in French, does it map these to the same internal representation of the feline concept, or does it maintain separate language-specific representations? Several recent works hint at the existence of language-agnostic concept representations.
@@
Summary of contributions. More specifically, in this work we aim to understand how transformers process and represent concepts during translation tasks, whether language and concept information can be manipulated independently in the model’s computations, and, whether models maintain separate language-specific concept representations or develop a shared conceptual space. To this end, we make the following contributions.
@@ 2 Related Work
LLMs have demonstrated remarkable capabilities in processing multilingual text across languages, with examples including encoder-only model like mBERT Devlin et al. (2018), XLM-R Conneau et al. (2020a), and mT5 Xue et al. (2021) and decoder-only model like Aryabumi et al. (2024); Dubey et al. (2024). Studies on encoder-only models have shown that they develop language-agnostic representations, explaining their cross-lingual transfer capabilities. The methodology used was embedding similarity analysis Wu et al. (2019); Libovický et al. (2020); Muller et al. (2021); Conneau et al. (2020b) and probing methods Choenni and Shutova (2020); Pires et al. (2019).
@@
Whiled decoder-only transformers are not primarily designed to develop contextual embedding, but rather for open-ended text generation, they also develop cross-lingual generalization, for example, when safety and instruction tuning Li et al. (2024); Chirkova and Nikoulina (2024). Mechanistic interpretability has led to powerful tools to analyze the language-agnosticity of these models. Using neuron analysis Stańczak et al. (2022); Chen et al. (2024b); Cao et al. (2024); Zeng et al. (2024); Tang et al. (2024) have shown that LLMs develop both language-agnostic and language-specific neurons. Wendler et al. (2024); Wu et al. (2024) use the logit lens Nostalgebraist (2020) to perform early decoding during the forward pass of both LLMs and VLMs and show that, no matter the language or input modality, the intermediate decodings concentrate on English before decoding to a specific language in the very last layers.
@@
While the analyses via embedding similarity, probing and the logit lens give use valuable insight into the structure of the representation, they are not causal. For example, using logit lens insights from Wendler et al. (2024); Wu et al. (2024) does not allow to do counter factual intervention Anonymous (2024). Additionally, while the neuron level analysis studies the causal effects of the neurons, they do not study the representation themselves. Our work aims to fill this gap.
@@
In order to do so, we draw inspiration from activation patching, which was introduced by Meng et al. (2022) and has been the main tool used to draw causal interpretation of LLM representations  Variengien and Winsor (2023); Geiger et al. (2022); Kramár et al. (2024). More recently, Ghandeharioun et al. (2024); Chen et al. (2024a) also show patching setups in which they can use the model itself to analyze its own model internal. Inspired by those methods, we developped two novel patching experiments supporting the language-agnosicity of LLMs representation.
@@ 3 Background
Transformers’s forward pass. When an autoregressive decoder-only transformer Vaswani et al. (2017); Touvron et al. (2023) with $L$ layers processes a sequence of input tokens $x_{1},\dots,x_{n}\in V$ from a vocabulary $V$, each token is initially transformed into a $d$-dimensional vector $h_{i}^{0}$ by an embedding layer. This first set of vector is the beginning of the residual stream. Then, for each token position $i$, the layer $j\in 1,\dots,L$ updates the residual stream the following way:
@@
Activation patching. Activation patching is a causal intervention technique that allows us to understand how different components of a neural network contribute to its output. The key idea is to run two forward passes through the network – one on a source input and one on a target input – and copy (or "patch") activations from specific positions and layers of the source forward pass into the target forward pass. By observing how these interventions affect the model’s output, we can understand what information is encoded in different parts of the network and how it is used.
@@
Prompt design. We use the same template as Wendler et al. (2024): For a given concept $C$, input language $\ell^{(\text{in})}$, and output language $\ell^{(\text{out})}$, we construct a few-shot translation prompt $\text{TP}(\ell^{(\text{in})},\ell^{(\text{out})},C)$. This prompt contains examples of single-word translations from $\ell^{(\text{in})}$ to $\ell^{(\text{out})}$, concluding with the model being tasked to translate $C$ from $\ell^{(\text{in})}$ to $\ell^{(\text{out})}$.
$\ell^{(\text{in})}$: “$C_{1}^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “$C_{1}^{\ell^{(\text{out})}}$ $\ell^{(\text{in})}$: “$C_{n}^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “$C_{n}^{\ell^{(\text{out})}}$ $\ell^{(\text{in})}$: “$C^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “
We denote $\text{TP}^{\text{collect}}(\ell^{(\text{in})},\ell^{(\text{out})},C)$ the translation prompt ending at the last token of $C^{\ell^{(\text{in})}}$.
@@
Importantly, whether the model correctly answers the prompt is determined by its next token prediction. For example, above, the next token predicted should be “nu", the first token of “nuage". Thus, we can track $P(C^{\ell})$, i.e., the probability of the concept $C$ occurring in language $\ell$, by simply summing up the probabilities of all starting tokens of $w(C^{\ell})$ in the next-token distribution.
@@ 4.1 Experimental design
Figure 2: For two given concepts, e.g., book and lemon, we construct a source prompt for translating book from German to Italian, and a target prompt for translating lemon from French to Chinese. Then we extract the residual stream of the last token after some layer $j$ from the source prompt and insert it at the corresponding position in the forward pass of the target prompt. The resulting next token probabilities will concentrate on the target concept in the target language ($\text{{{lemon}}}^{\text{{{zh}}}}$, i.e., {CJK}UTF8gbsn柠檬) when patching at layers 0–11, on the target concept in the source language ($\text{{{lemon}}}^{\text{{{it}}}}$, “limone”) for layers 12–16, and on the source concept in the source language ($\text{{{book}}}^{\text{{{it}}}}$, “libro”) for layers 17–31.
We start our analysis with an exploratory experiment on Llama 2 7B Touvron et al. (2023). We use 5-shots translation prompts to create paired source $S=\text{TP}(\ell^{(\text{in})}_{S},\ell^{(\text{out})}_{S},C_{S})$ and target prompt $T=\text{TP}(\ell^{(\text{in})}_{T},\ell^{(\text{out})}_{T},C_{T})$ datasets with different concept, input languages and output languages. If not mentioned otherwise, $\ell_{S}$ and $\ell_{T}$ refer to the output language of $S$ and $T$.
Similar to Variengien and Winsor (2023), we would like to infer at which layers the output language and the concept enter the residual stream $h_{n_{T}}^{(j)}(T)$ respectively and whether they can vary independently for our task. In order to investigate this question, we perform the experiment depicted in Figure 2. For each transformer block $f_{j}$ we create two parallel forward passes, one processing the source prompt $S=(s_{1},\dots,s_{n_{S}})$ and one processing the target prompt $T=(t_{1},\dots,t_{n_{T}})$. While doing so, we extract the residual stream of the last token of the source prompt after layer $j$, $h_{n_{S}}^{(j)}(S)$, and insert it at the same layer at position $n_{T}$ in the forward pass of the target prompt, i.e., by setting $h_{n_{T}}^{(j)}(T)=h_{n_{S}}^{(j)}(S)$ and subsequently completing the altered forward pass. From the resulting next token distribution, we compute $P(C_{S}^{\ell_{S}}),P(C_{S}^{\ell_{T}}),P(C_{T}^{\ell_{S}}),\text{ and }P(C_{T}^{\ell_{T}})$.
@@ 4.2 Results
Figure 3: Our first patching experiment with a de to it source prompt and a fr to zh target prompt with different concepts. The x-axis shows at which layer the patching was performed and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). In the legend, the prefix “src" stands for source and “tgt" for target concept. The orange dashed line and blue dash-dotted line correspond to the mean accuracy on source and target prompt. We report means and 95% Gaussian confidence intervals computed over 200 source, target prompt pairs featuring 41 source concepts and 38 target concepts.
@@
In later layers, the model determines the concept: from layer 16 onwards, the source concept is decoded. This suggests that $z_{C_{T}^{\ell^{(\text{out})}}}$ is overwritten at layer 16.222In Appendix A, we collected additional experimental results investigating the right part of Figure 3 more deeply and in Appendix B the left part. For the right part, we use the patchscope lens Ghandeharioun et al. (2024) to investigate from which layer it is possible to decode the source concept in App. Figure 7. The results of both experiments agree: from layer 16 it is possible to decode the source concept in source language. For the left part, we experiment with randomized source prompts and different prompting templates in between source and target prompt in App. Figure 8. We find that indeed before layer 11 there is no translation task specific information in the residual stream, only prompt-template specific information.
@@ 5 Ruling out hypotheses
Further evidence experiment. In the experiments in Sec. 4 we did not observe source concept in target language. However, both H1 and H2 would allow for that to happen via patching in the right way. Therefore, in this experiment, instead of overwriting the residual stream at the last token of the prompt, we overwrite them at the last token of the word to be translated. In order to do that, for the source prompt, we use $\text{TP}^{\text{collect}}$ instead of TP ($S=\text{TP}^{\text{collect}}(\ell^{(\text{in})}_{S},\ell^{(\text{out})}_{S},C_{S})$). This means that we collect the activations at the last token of $C_{S}^{\ell_{S}^{(\text{in})}}$.
@@
Formally, we patch by setting $h_{\rho_{T}}^{(j)}(T)=h_{-1}^{(j)}(S),\dots,h_{\rho_{T}}^{(L)}(T)=h_{-1}^{(L)}(S)$.
@@
Therefore, we’d expect high $P(C_{S}^{\ell_{T}})$ in this case. However, under H2, in which $z_{C_{i}}$ cannot be disentangled, this mean may not correspond to a well-defined concept. Additionally, the interference between multiple input languages should cause difficulties for the language-pair-specific translation, which should result in a drop in $P(C_{S}^{\ell_{T}})$.
@@
Figure 4: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use multiple source input languages de, nl, zh, es, ru and source output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt in (b) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src" stands for source and “tgt" for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
Other models. In Appendix C we perform the experiments from Sec. 4 and Sec. 5 on several other models, varying in size, training data and architecture namely, Mistral 7B Jiang et al. (2023), Llama 3 8B Dubey et al. (2024), Qwen 1.5 7B Bai et al. (2023), Llama 2 70B, and Aya 23 8B Aryabumi et al. (2024) which was specifically trained to be multilingual. We observe the same improvement when we take the mean of a concept across languages for all these models, suggesting that they are all operating under H1 and use some language-agnostic concept representation.

### v4

@@ Abstract
A central question in multilingual language modeling is whether large language models (LLMs) develop a universal concept representation, disentangled from specific languages. In this paper, we address this question by analyzing latent representations (latents) during a word-translation task in transformer-based LLMs. We strategically extract latents from a source translation prompt and insert them into the forward pass on a target translation prompt. By doing so, we find that the output language is encoded in the latent at an earlier layer than the concept to be translated. Building on this insight, we conduct two key experiments. First, we demonstrate that we can change the concept without changing the language and vice versa through activation patching alone. Second, we show that patching with the mean representation of a concept across different languages does not affect the models’ ability to translate it, but instead improves it. Finally, we generalize to multi-token generation and demonstrate that the model can generate natural language description of those mean representations. Our results provide evidence for the existence of language-agnostic concept representations within the investigated models. [^1] This work has been previously published under the title “How Do Llamas Process Multilingual Text? A Latent Exploration through Activation Patching” at the ICML 2024 Mechanistic Interpretability Workshop https://openreview.net/forum?id=0ku2hIm4BS.
Code for reproducing our experiments is available at https://github.com/Butanium/llm-lang-agnostic
@@
Clément Dumas12thanks: Equal contributionthanks: Work done while at EPFL  Chris Wendler311footnotemark: 122footnotemark: 2 Veniamin Veselovsky422footnotemark: 2  Giovanni Monea522footnotemark: 2  Robert West6 1ENS Paris-Saclay  2Université Paris-Saclay  3Northeastern  4 Princeton  5Cornell  6EPFL {clement.dumas@ens-paris-saclay.fr, chris.wendler@epfl.ch}
@@ 1 Introduction
A fundamental question in multilingual language modeling is whether LLMs develop universal concept representations that transcend specific languages Wendler et al. (2024); Conneau et al. (2020b); Chi et al. (2020); Xie et al. (2022); Mousi et al. (2024). For example, when a model processes the word “cat” in English and “chat” in French, does it map these to the same internal representation of the feline concept, or does it maintain separate language-specific representations? Several recent works hint at the existence of language-agnostic concept representations.
@@
Contributions. While prior work has provided observational evidence for shared semantic spaces in LLMs, we present the first causal analysis of how these representations are actively utilized during multilingual processing. More specifically, in this work we aim to understand how transformers process and represent concepts during translation tasks, whether language and concept information can be manipulated independently in the model’s computations, and, whether models maintain separate language-specific concept representations or develop a shared conceptual space. To this end, we make the following contributions.
@@
Implications. While prior work has suggested the existence of shared semantic spaces across languages through observational methods, our causal analysis provides the first direct evidence that LLMs actively utilize language-agnostic concept representations during text generation. Furthermore, our activation patching methodology establishes a framework for future causal investigations of multilingual representations, moving beyond the limitations of embedding similarity, probing, and logit lens approaches.
@@ 2 Related Work
LLMs have demonstrated remarkable capabilities in processing multilingual text across languages, with examples including encoder-only model like mBERT Devlin et al. (2018), XLM-R Conneau et al. (2020a), and mT5 Xue et al. (2021) and decoder-only model like Aryabumi et al. (2024); Dubey et al. (2024). Studies on encoder-only models have shown that they develop language-agnostic representations, explaining their cross-lingual transfer capabilities. The methodology used was embedding similarity analysis Conneau et al. (2020b); Libovický et al. (2020); Muller et al. (2021); Mousi et al. (2024) and probing methods Choenni and Shutova (2020); Pires et al. (2019).
@@
While decoder-only transformers are not primarily designed to develop contextual embedding, but rather for open-ended text generation, they also develop cross-lingual generalization, for example, during safety and instruction tuning Li et al. (2024); Chirkova and Nikoulina (2024). Mechanistic interpretability has led to powerful tools to analyze the language-agnosticity of these models. Using neuron analysis Stańczak et al. (2022); Chen et al. (2024b); Cao et al. (2024); Zeng et al. (2024); Tang et al. (2024) have shown that LLMs develop both language-agnostic and language-specific neurons. Wendler et al. (2024); Wu et al. (2024) use the logit lens Nostalgebraist (2020) to perform early decoding during the forward pass of both LLMs and show that, no matter the language or input modality, the intermediate decodings concentrate on English before decoding to a specific language in the very last layers.
@@
While the analyses via embedding similarity, probing and the logit lens give use valuable insight into the structure of the representation, they are not causal. Additionally, while the neuron level analysis studies the causal effects of the neurons, they do not study the representation themselves. Our work aims to fill this gap.
@@
A related line of work has explored using definition generation as a means to evaluate semantic representations. Noraset et al. (2017) introduced definition modeling, the task of generating dictionary definitions from word embeddings, as a more direct evaluation of what semantic information embeddings capture. This approach has been extended to evaluate various types of embeddings Gardner et al. (2022); Chang and Chen (2019). For example, Chang and Chen (2019) demonstrated that contextualized embeddings like ELMo and BERT can be effectively mapped to definition spaces, revealing their sense-specific semantic content. In our work, instead of training a model to generate descriptions of LLM’s representations, we repurpose the LLM itself to either translate or define it using activation patching. Activation patching, introduced by Meng et al. (2022), has been the main tool used to draw causal interpretation of LLM representations  Variengien and Winsor (2023); Geiger et al. (2022); Kramár et al. (2024). More recently, Ghandeharioun et al. (2024); Chen et al. (2024a) also show patching setups in which they can use the model itself to analyze its own model internal. Inspired by those methods, we developed two novel patching experiments supporting the language-agnosicity of LLMs representation.
Parallel work to ours by Fierro et al. (2025) leverages the mechanistic interpretability toolkit to understand how factual recall works in multilingual LLMs as well. Similar to us, they find that the concept to be decoded and its language does not enter at the same layer into the residual stream, however in their knowledge association tasks they observe the opposite order than we do in our translation task. They also find that large parts of multilingual-factual recall are handled in a language agnostic way, despite the investigated models being trained on a more balanced split of languages than the ones studied in our paper.
@@ 3 Background
Transformers forward pass. When an autoregressive decoder-only transformer Vaswani et al. (2017); Touvron et al. (2023) with $L$ layers processes a sequence of input tokens $x_{1},\dots,x_{n}\in V$ from a vocabulary $V$, each token is initially transformed into a $d$-dimensional vector $h_{i}^{0}$ by an embedding layer. This first set of vector is the beginning of the residual stream. Then, for each token position $i$, the layer $j\in 1,\dots,L$ updates the residual stream the following way:
@@
Activation patching. Activation patching is a causal intervention technique that allows us to understand how different components of a neural network contribute to its output. The key idea is to run two forward passes through the network – one on a source input and one on a target input – and copy (or “patch”) activations from specific positions and layers of the source forward pass into the target forward pass. By observing how these interventions affect the model’s output, we can understand what information is encoded in different parts of the network and how it is used.
@@
More formally, for a given concept $C$, input language $\ell^{(\text{in})}$, and output language $\ell^{(\text{out})}$, we construct a few-shot translation prompt $\text{TP}(\ell^{(\text{in})},\ell^{(\text{out})},C)$. This prompt contains $k$ examples222in our study we used $k=5$ of single-word translations of concepts $C_{1},\dots,C_{k}$ from $\ell^{(\text{in})}$ to $\ell^{(\text{out})}$, concluding with the model being tasked to translate $C$ from $\ell^{(\text{in})}$ to $\ell^{(\text{out})}$. Using $C^{\ell^{(\text{in})}}$ as a shortcut for $v\in w(C^{\ell^{(\text{in})}})$, $\text{TP}(\ell^{(\text{in})},\ell^{(\text{out})},C)$ looks like:
$\ell^{(\text{in})}$: “$C_{1}^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “$C_{1}^{\ell^{(\text{out})}}$ $\ell^{(\text{in})}$: “$C_{k}^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “$C_{k}^{\ell^{(\text{out})}}$ $\ell^{(\text{in})}$: “$C^{\ell^{(\text{in})}}$” - $\ell^{(\text{out})}$: “
We denote $\text{TP}^{\text{concept}}(\ell^{(\text{in})},\ell^{(\text{out})},C)$ as $\text{TP}(\ell^{(\text{in})},\ell^{(\text{out})},C)$ cut at the last token of $C^{\ell^{(\text{in})}}$. For example, in our previous example, $\text{TP}^{\text{concept}}(\textsc{{en}},\textsc{{fr}},\textsc{{cloud}})$ would be:
English: “computer” - Français: “ordinateur” English: “ant” - Français: “fourmi” English: “cloud
Importantly, whether the model correctly answers TP is determined by its next token prediction. In our prompt example, the next token predicted should be “nu”, the first token of “nuage”. Thus, we can track $P(C^{\ell})$ [^3] We use simplified notation $P(C^{\ell})$ rather than $P(C^{\ell}|\text{TP})$ throughout. While the conditional notation would be more precise for the initial case, our patching experiments involve multiple conditioning factors (target prompt, source prompt, patching configuration, layer) that would make the full notation unwieldy. We therefore adopt this simplified notation for clarity., i.e., the probability of the concept $C$ occurring in language $\ell$, by simply summing up the probabilities of all starting tokens of $w(C^{\ell})$ in the next-token distribution.
@@ 4.1 Experimental design
Figure 2: For two given concepts, e.g., book and lemon, we construct a source prompt for translating book from German to Italian, and a target prompt for translating lemon from French to Chinese. Then we extract the residual stream of the last token at a single layer $j$ from the source prompt and insert it at the corresponding position and layer in the forward pass of the target prompt. The resulting next token probabilities will concentrate on the target concept in the target language ($\text{{{lemon}}}^{\text{{{zh}}}}$, i.e., {CJK}UTF8gbsn柠檬) when patching at layers 0–11, on the target concept in the source language ($\text{{{lemon}}}^{\text{{{it}}}}$, “limone”) for layers 12–16, and on the source concept in the source language ($\text{{{book}}}^{\text{{{it}}}}$, “libro”) for layers 17–31.
We start our analysis with an exploratory experiment on Llama 2 7B Touvron et al. (2023). We use 5-shots translation prompts to create paired source $S=\text{TP}(\ell^{(\text{in})}_{S},\ell^{(\text{out})}_{S},C_{S})$ and target prompt $T=\text{TP}(\ell^{(\text{in})}_{T},\ell^{(\text{out})}_{T},C_{T})$ datasets with different concept, input languages and output languages [^4] See details in Appendix B. If not mentioned otherwise, $\ell_{S}$ and $\ell_{T}$ refer to the output language of $S$ and $T$.
Similar to Variengien and Winsor (2023), we would like to infer at which layers the output language and the concept enter the residual stream $h_{n_{T}}^{(j)}(T)$ respectively and whether they can vary independently for our task. In order to investigate this question, we perform the experiment depicted in Figure 2. For each transformer block $f_{j}$ we create two parallel forward passes, one processing the source prompt $S$ which tokens are $(s_{1},\dots,s_{n_{S}})$ and one processing the target prompt $T=(t_{1},\dots,t_{n_{T}})$. While doing so, we extract the residual stream of the last token of the source prompt after layer $j$, $h_{n_{S}}^{(j)}(S)$, and insert it at the same layer at position $n_{T}$ in the forward pass of the target prompt, i.e., by setting $h_{n_{T}}^{(j)}(T)=h_{n_{S}}^{(j)}(S)$ and subsequently completing the altered forward pass. From the resulting next token distribution, we compute $P(C_{S}^{\ell_{S}}),P(C_{S}^{\ell_{T}}),P(C_{T}^{\ell_{S}}),\text{ and }P(C_{T}^{\ell_{T}})$.
@@ 4.2 Results
Figure 3: Our first patching experiment with a de to it source prompt and a fr to zh target prompt with different concepts. The x-axis shows at which layer the patching was performed and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). In the legend, the prefix “src” stands for source and “tgt” for target concept. The orange dashed line and blue dash-dotted line correspond to the mean accuracy on source and target prompt. We report means and 95% Gaussian confidence intervals computed over 200 source, target prompt pairs featuring 41 source concepts and 38 target concepts.
@@
In later layers, the model determines the concept: from layer 16 onwards, the source concept is decoded. This suggests that $z_{C_{T}^{\ell^{(\text{out})}}}$ is overwritten at layer 16.555In Appendix A, we collected additional experimental results investigating the right part of Figure 3 more deeply and in Appendix C the left part. For the right part, we use the patchscope lens Ghandeharioun et al. (2024) to investigate from which layer it is possible to decode the source concept in App. Figure 7. The results of both experiments agree: from layer 16 it is possible to decode the source concept in source language. For the left part, we experiment with randomized source prompts and different prompting templates in between source and target prompt in App. Figure 8. We find that indeed before layer 11 there is no translation task specific information in the residual stream, only prompt-template specific information.
@@ 5 Ruling out hypotheses
Further evidence experiment. In the experiments in Sec. 4 we did not observe source concept in target language. However, both H1 and H2 would allow for that to happen via patching in the right way. Therefore, in this experiment, instead of overwriting the residual stream at the last token of the prompt, we overwrite them at the last token of the word to be translated. In order to do that, for the source prompt, we use $\text{TP}^{\text{concept}}$ instead of TP ($S=\text{TP}^{\text{concept}}(\ell^{(\text{in})}_{S},\ell^{(\text{out})}_{S},C_{S})$). This means that we collect the activations at the last token of $C_{S}^{\ell_{S}^{(\text{in})}}$.
@@
Formally, we patch by setting $h_{\rho_{T}}^{(j)}(T)=h_{-1}^{(j)}(S),\dots,h_{\rho_{T}}^{(L)}(T)=h_{-1}^{(L)}(S)$ [^6] Note that we use Python indexing, where -1 denotes the last token..
@@
Therefore, we’d expect high $P(C_{S}^{\ell_{T}})$ in this case. However, under H2, in which $z_{C_{i}}$ cannot be disentangled, this mean representation may not correspond to a well-defined concept. Additionally, the interference between multiple input languages should cause difficulties for the language-pair-specific translation, which should result in a drop in $P(C_{S}^{\ell_{T}})$.
@@
Figure 4: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use multiple source input languages de, nl, zh, es, ru and source output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt in (b) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src” stands for source and “tgt” for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
Other models. In Appendix D we perform the experiments from Sec. 4 and Sec. 5 on several other models, varying in size, training data and architecture namely, Mistral 7B Jiang et al. (2023), Llama 3 8B Dubey et al. (2024), Qwen 1.5 7B Bai et al. (2023), Llama 2 70B, and Aya 23 8B Aryabumi et al. (2024) which was specifically trained to be multilingual. We observe the same improvement when we take the mean of a concept across languages for all these models, suggesting that they are all operating under H1 and use some language-agnostic concept representation.